Frame Masters, Inc., recently purchased $5500 worth of new equipment in order to offer a new eyeglass frame. The marketing department of Frame Masters estimates that the cost of producing this new fame is $18 and that the frame will be sold to stores for $38. Find the number if frames that must be sold in order to break even.
Answer by Theo(13342)

the equipment to make the frames is equal to 5500.
the cost of producing the new frames is 18 dollars each.
this is presumably on top of the cost of the equipment used.
the frames will be sold for 38 dollars each.

break even point is when revenue equals cost.

x = number of frames sold

cost = 5500 + 18x

revenue = 38x

revenue = cost when 38x = 5500 + 18x

subtract 18x from both sides of this equation to get 20x = 5500

solve for x to get x = 5500/20 = 275

they need to sell 275 eyeglass frames to break even.

38*275 = 10450
5500 + 18*275 = 10450